Swiss Furniture Company manufactures bookshelves and uses an activity−based costing system to allocate all manufacturing conversion costs. The company incurs no direct labor costs. The following information is provided for the month of​ May:

Activity

Estimated Indirect

Activity Costs

Allocation Base

Estimated Quantity of

Allocation Base

Materials handling

​$7,500

Number of parts

​7,080 parts

Assembling

​$11,900

Number of parts

​7,080 parts

Packaging

​$2,160

Number of bookshelves

​1,180 bookshelves

Each bookshelf consists of 6 parts. The direct materials cost per bookshelf is​ $35.

What is the total manufacturing cost per​ bookshelf?

A.

​$18.27

B.

​$16.44

C.

​$51.44

D.

​$53.27
Answer #1

Calculation of manufacturing cost per bookshelf:

Direct Material =$35

Overheads = Cost*Activity consumed/Total Activity

Material Handling = 7,500*6/7,080 = $6.36

Assembling = 11,900*6/7,080 = $10.08

Packaging = 2,160*1/1,180 = $1.83

Total manufacturing cost per bookshelf = $53.27

Hence, the answer is D.